OpenStudy (anonymous):

ben is making waffles for his 3 friends .the recipe calls for 2/3 cup of bleached flour and 1/4 cup of wheat flour .how many cups of wheat flour does been neeed .?

OpenStudy (anonymous):

do you give enough information? question incomplete?? pls check??

OpenStudy (anonymous):

thats it

OpenStudy (anonymous):

i hope that, answer is 3* 1/4 = 3/4

OpenStudy (anonymous):

3/4 cups of wheat flour for making waffles for his 3 friends

OpenStudy (anonymous):

should i add the denominators

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for the case of bleached flour , answer will be 3*2/3 = 2 cup but for wheat answer is 3*1/4 = 3/4 cup clear?/

OpenStudy (anonymous):

or do i subtract them